OpenStudy (anonymous):

Can someone explain this to me, step by step: Simplify completely quantity 4x+16 over 6.

OpenStudy (alexandervonhumboldt2):

\[\frac{ 4x+16 }{ 6 }=\frac{ 4*(x+4) }{ 2*3 }=\frac{ 4 }{ 2 }*\frac{ x+4 }{ 3 }=\frac{ 2*(x+4) }{ 3 }=distribute ~~and ~~simplify\]